CAT Chat Minutes

April 3, 1998

Issues that require action and/or follow-up

1. Will there be a discussion on the long-term schedule (starting in Oct. '98) today?

We are still working on the plan for tunnel access to make repairs. We are not ready at this time to present this information. We have the beginning and the end dates of the last cycle in 1998 and will get that information out to you via a CAT net.

2. Everything in TB 20 refers to the ID. Is there a plan to write a bulletin that deals with the BM ?

Yes, we will do this. P.K. Job has all the necessary information.

3. When the station door has closed and locked, why is there a 20 second waiting period on the PSS system? Can this be changed?

The 20 second delay is designed to allow anyone locked inside the station adequate time to hit the emergency shutdown button. It is a requirement for engineered safety systems.

4. Is there any update on the liquid nitrogen systems?

We have been told by DOE, that the proposal from APS CATs will not be funded in the second half of this fiscal year. No assurances for funding for the next fiscal year, either. We will continue to make this request through the APS CAT committee.

5. On 5 ID B, C, and D, we noticed that the EPS light shows green. Since everything is off-line, this is clearly not the case. The light shows red on the A Station. What does this mean? It is confusing for our users.

We will check this out, clarify it for you, and take care of this.

Information/Follow-up Items

APS Operational Status Update

Users Operation continues until 0800 on Monday, April 6, followed by 24 hours of Machine Studies until 0800 on Tuesday, April 7. At that time, we have a shift of shielding verification at 10 and 25 mA. We may not require the full shift. Beginning at 1600 hours Tuesday, we will start the SOM of singlets. We will be tweaking this mode to allow us to be able to fill to 100 mA. We plan on two fills per day.

Information

The ring road lane painting from parking lots and stop signs at the 431 truck lock area will begin the week of April 13.

The FBI will be working here. Their base of operation will be the 402 Gallery, but most of their activities will be in other parts of the lab.

A couple of weeks back, a request was made to have maps put up in this building directing visitors to areas such as the stockroom, etc. The maps are being printed and will be posted soon.

B. Ferry in now in charge of all floor operations. We recently made a change, giving Bob the responsibility for this area.

We are still collecting information on 5 ID. We do not at this time have a handle on what caused this problem. We can see where the beam has hit. There is a group of people looking into this issue and writing the procedure on dismantling the collimator. We will keep you informed.

Safety

In another area of the lab, there was an incident with the operation of a class 4 laser that resulted in some retina damage. This is just a reminder to be cautious when using lasers.